Consider a one dimensional monatomic lattice with lattice constant 4.91A and atoms of mass 28amu. The speed of sound in the lattice is 11.6kms-1 . A) calculate the maximum frequency of waves that can be transmitted by the lattice.B) calculate the interatomic force constant of the lattice
1
Expert's answer
2020-04-13T10:21:23-0400

Molecular Physics | Thermodynamics

We need to calculate the maximum frequency of waves and calculate the inter atomic force

Solution:

Given

Mass of atoms = "m = 28 \space amu = 28 \times 1.67 \times 10^{-27} kg"


speed of sound in the lattice ="v_s= 11.6 \space km /s = 11.6 \times 10^3 m/s"


lattice constant = "a = 4.91A^o = 4.91 \times 10^{-10} m/s"


A).


 The maximum frequency of waves which was transmitted by the lattice is

"w_m =\frac {2}{a} \times v_s = \frac {2}{4.91 \times 10^{-10} } \times 11.6 \times 10^3"

"w_m =4.725 \times 10^{13} rad/sec"

B).

"w_m = \sqrt {\frac {4\beta}{m}}"

  Squaring on both sides

"(w_m)^2 = {\frac {4\beta}{m}}"

"\beta = \frac {m \times \beta^2 } {4} = \frac {28 \times (4.725 \times 10^{13})^2 } {4}"

"\beta = \frac {28 \times 1.67 \times 10^{-27} \times 22.325 \times 10^{26}} {4}"

"\beta = \frac {1043.917 \times 10^{-1}}{4} =26.0979 N/m"


Answer: "\beta = 26.0979 \space N/m"